OpenStudy (firejay5):

@Mertsj and @satellite73 Simplify each expression. 16. 3x + 6 over x^2 - 9 divided by 6x^2 + 12x over 4x + 12

OpenStudy (anonymous):

\[\frac{3x+6}{x^2-9}\times \frac{4x+12}{6x^2+12x}\] is a start, then factor a bunch and cancel a bunch

OpenStudy (firejay5):

Can you show me, because I am totally stuck on this problem

OpenStudy (anonymous):

first step is to factor \[\frac{3(x+2)4(x+3)}{(x+3)(x-3)6x(x+2)}\] then get rid of the common factors

OpenStudy (firejay5):

Cause the next part confused me

OpenStudy (anonymous):

then you are done

OpenStudy (firejay5):

What would it look like when it's done

OpenStudy (anonymous):

you should only be left with \(\frac{2}{x-3}\)

OpenStudy (firejay5):

How did you get that?

OpenStudy (anonymous):

\[\frac{3\cancel{(x+2)}4(x+3)}{(x+3)(x-3)6x\cancel{(x+2)}}\] \[\frac{12(x+3)}{6x(x+3)(x-3)}\] \[\frac{12\cancel{(x+3)}}{6x\cancel{(x+3)}(x-3)}\]\[\frac{12}{6x(x-3)}\]

OpenStudy (anonymous):

then finally cancel the 6

OpenStudy (firejay5):

so the final answer would be: 2 over x - 3

OpenStudy (mertsj):

\[\frac{2}{x(x-3)}\]

OpenStudy (mertsj):

That's the correct answer.

OpenStudy (firejay5):

why is the 2 on top

OpenStudy (mertsj):

Because 12/6 = 2

OpenStudy (firejay5):

because 12 is the bigger number right

OpenStudy (mertsj):